OpenStudy (anonymous):

Lonnie has a gift card worth $600 for a local entertainment store. Movies cost $20 each and newly released video games cost $50 each. In order to use the gift card, Lonnie must purchase at least thirteen items. If Lonnie wants to spend as much as possible on his gift card by purchasing the most items, which combination of movies and newly released video games is optimal? Help me please ^_^

OpenStudy (anonymous):

@iGreen HELP ME OUT PLZZZ

OpenStudy (igreen):

Okay, this is similar to the last one.

OpenStudy (igreen):

Our first equation is 20x + 50y = 600. Where x = movies, and y = video games. Our second equation is x + y = 13.

OpenStudy (anonymous):

i GOT (10, 8)

OpenStudy (igreen):

If you graph it you get \((\dfrac{1}{6}, 11 \dfrac{1}{3})\)

OpenStudy (anonymous):

SO ITS (20, 4)